<e> Examine the f:R->R defined by



f(x)={1/6(x+1)^3 x is not equal to 0



5/6 x=0}



for continuity on R .If it is not continuous at any of R find the nature of discontinuity there



1
Expert's answer
2022-02-16T10:19:27-0500

Since


"\lim\limits_{x\to 0}f(x)=\lim\limits_{x\to 0}\frac{1}6(x+1)^3=\frac{1}6\ne\frac{5}6=f(0),"


we conclude that "x=0" is a point of removable discontinuity.


On the set "\R\setminus\{0\}" the function "f" is an elementary function, and hence it is continuous.